OpenStudy (anonymous):

solve by substitution 2x-y=7 y=2x-5

OpenStudy (anonymous):

ok substitute 2x-(2x-5)=7

OpenStudy (anonymous):

OK IT CAN BE SOLVED BY COLUMN METHOD

OpenStudy (anonymous):

Its already solved for y in the 2nd equation 2x-(2x-5)=7 next combine like terms and divide by the coefficient 2x-2x+5=7 the 2x's cancel and \[5\neq7\]

OpenStudy (anonymous):

SEE HERE : 2x-y=7 and y=2x-5 let 2x-y = 7 equation be (1) and y=2x-5 equation be (2) now in the equation 2 we will transpose 2x to LHS or subtract both sides by 2x we would get: y-2x= -5 or 2x-y=5 2x-y=7 and 2x-y=5 let us subtract both equations (1) and (2) 2x-y-2x+y=7-5 0 \[\neq\]2 hence the given equations are wrong

OpenStudy (anonymous):

HEY GIVE ME MEDAL